The latent heat of fusion of ice value in C.G.S system in 80 cal/g. its value in SI system. Take 1 cal = 4.2 J

Grade:12th pass

2 Answers

User
7 years ago
In SI the value of latest heat of fusion of ice is given by80*4.2= 336 joules/g But the correct value of 1 cal is equal to 4.18 so after this you will calculate the answer would by as follow80*4.18= 334.4
Ahana Mukherjee
6 years ago
Approx value=80*4.2=336J/gReal value=80*4.18=334.4Other than this if I want any perfect answer,refer Google